OpenStudy (anonymous):

Two sides of right triangl are 6 and 8, what is a posssible value of the 3rd side? Choices are 15 11 5 rad 3 2 rad 7 2 rad 3 Note need explanation please

OpenStudy (hoblos):

8^2 = 6^2 + x^2 x= sqrt(64-36) = sqrt(28) = 2 sqrt(7)

OpenStudy (anonymous):

use pythagorean theorem just how hoblos explained

OpenStudy (anonymous):

3,4,5 is a Pythagorean triple

OpenStudy (anonymous):

and so is 6,8,10

OpenStudy (hoblos):

you are right ffm... but 10 is not one of the options

OpenStudy (anonymous):

Great job Hobls - yeah we were on that path and my son said that you are a Genius. Thanks guys.

OpenStudy (anonymous):

Hm yes, then we meed to assume 8 to be the hypotenuse.

OpenStudy (anonymous):

Apparently, we don't need to be a genius to answer this question ;-)

OpenStudy (hoblos):

lol... @FoolForMath is right

OpenStudy (phi):

It immediately think 3,4,5 --> 6,8,10 but no 10! so 8 is the hypotenuse (longest side) so sqrt(64-36) = 2sqrt(7)

hero (hero):

|a - b|<c<|b-a|

OpenStudy (anonymous):

i did that hero and there aree 4 answers that satisfy that - helping my son study for sat

hero (hero):

You should have known to just use pythagorean theorem

OpenStudy (anonymous):

right, makes sense now

hero (hero):

I have a feeling your son will do great on the SAT

hero (hero):

@JuanitaM

OpenStudy (anonymous):

Thanks Hero - he enjoyed your answer. This is his first time taking it so I would think he would want to score in his range to be done. UGH.... I def do not misss those days.

hero (hero):

I haven't taken the SAT yet